OpenStudy (anonymous):

Determine the solutions to the quadratic equation -5x2 + 10x + 15 = 0.

OpenStudy (anonymous):

You the quadratic formula or graph it

OpenStudy (anonymous):

*use

OpenStudy (wolf1728):

x = [-b +- sqrt(b² - 4 * a * c)] / 2 * a x = [-10 +- sqrt(100 - 4*-5*15)] / 2 * -5 x = [-10 +- sqrt (100 - -300)] / -10 x = [-10 +-sqrt(400)] / -10 x = [-10 +-20]/ -10 x = -10 + 20 / -10 = -1 x = -10 -20 / -10 = 3